# MRAG-Suite: A Diagnostic Evaluation Platform for Visual Retrieval-Augmented Generation

### v1

@@ 5.2 End-task Accuracy (short-answer EM/Acc)
Table 13 summarizes EM/Accuracy across datasets. On the Full split, LLaVA-2 reaches an overall accuracy of 62.5%. After applying our two-step filtering, accuracy drops to 47.8% (–14.7 points), confirming that the retained questions are substantially harder and more retrieval-dependent. Performance varies widely by domain: WebQA remains the easiest for the baseline (52% on the filtered set), while Chart-RAG is the hardest (low 30%s). Visual-RAG sits in between ( 45%), indicating persistent difficulty with fine-grained visual recognition. A no-retrieval ablation reduces accuracy to 25% on the filtered split, showing that at least half of the questions truly require external evidence. Manual inspection of 50 incorrect cases reveals two major categories: (a) Visual under-utilization (40%): the model retrieves the correct image/chart but misreads or ignores it; (b) Multi-hop failures (30%): the model retrieves partial evidence but fails to integrate it. The rest are simple reasoning mistakes or retriever misses.
@@ 5.3 Claim-level Diagnostics (MM-RAGChecker)
Using MM-RAGChecker, Table 3 summarizes claim-level diagnostics on the Filtered split. For LLaVA-2, 68% of claims are supported (hallucination 32%); ROUGE-L is 0.61. Faithfulness varies by domain, with WebQA higher ($\sim$72%) than Visual-RAG ($\sim$60%).
Hallucinations in Long Answers.
MM-RAGChecker shows that about one-third of answer statements are unsupported. Most are intrinsic hallucinations—plausible but unevidenced details (e.g., adding a material or breed not visible in the image)—while extrinsic hallucinations (off-topic filler) are rarer. For instance, in a Visual-RAG sample the model correctly states that a dog is catching a frisbee, but hallucinates its color and breed; our checker flags only those clauses.
Cross-modality behavior.
Beyond per-claim labels, we analyze modality interplay in Table 4, using $\DeltaCR/Δ\DeltaCP$ (image vs. text claim recall/precision gaps), VIS-Hit@$k$ and TXT-MissRate (evidence coverage), CMA (cross-modality agreement), and V-HR/D-HR (where hallucinations are mistakenly “entailment”). Large positive $\DeltaCR/Δ\DeltaCP$ indicates an image-leaning model, while high TXT-MissRate reveals wasted textual context. Lower CMA suggests the model tends to commit to a single modality instead of reconciling both.
Evidence Utilization and Modality Bias.
On average, the model uses $\sim$21% of retrieved pieces—suggesting poor integration. We observe a bias to stick to either text or image evidence instead of combining both, echoing prior findings that VLMs over-rely on text. Providing only gold evidence (GO mode) increases the Supported fraction to $\sim$90%, indicating distractors significantly hurt grounding.
Evidence Utilization.
Across datasets, LLaVA-2 uses on average 2.1 out of 10 retrieved items ($\sim$21% utilization), indicating under-use of available context. When provided only gold evidence (GO mode), the supported claim ratio climbs to about 90%, implying distractors directly harm grounding.
@@ 5.4 Effect of Retrieval Modes
Table 5 shows EM/Accuracy under DO, GPD, and GO. DO severely degrades performance (as expected), while GPD lies between DO and GO. The gap between GPD and GO quantifies the model’s susceptibility to distractors; Chart-RAG and VisRAG-Doc exhibit the largest drops, revealing brittle selection of evidence.
@@ 5.5 Prompt Sensitivity Study
We sweep 12 prompt designs (init style, few-shot exemplar count, reasoning pattern). Table 6 lists the best-performing prompt per dataset; Figure 3 visualizes the direct vs. retrieve-then-reason trend.
Table 6 reports, for each dataset, the prompt that achieves the highest mean Accuracy/EM across six models (filtered split). We observe clear patterns: Reasoning matters: retrieve_then_reason dominates on text-heavy retrieval tasks (MRAG, WebQA), whereas direct is consistently better on VisRAG-* visual QA splits where context is already focused. Few-shot count is task-dependent: Using three exemplars (ex3) suffices—and sometimes outperforms ex6—for purely visual/numeric QA (plots/slides). In contrast, ex6 helps more on document-style and open-web questions that benefit from pattern coverage. Init style is secondary: Switching between plain and expert seldom changes the rank unless combined with the right reasoning pattern and example count.
If a single prompt must be used across all datasets, expert-ex3-direct provides the best overall trade-off (top-1 on 3/6 datasets and close second elsewhere).
@@ 5.7 Effect of Filtering
Filtering not only reduces accuracy but also changes answer style. Average long-answer length increases from 45 to 60 words after filtering, yet factuality does not improve—MM-RAGChecker flags many of the additional sentences as unsupported. This suggests that harder questions trigger verbosity rather than better grounding.
5.8 Takeaways
(1) Filtering and distractors expose large drops in both accuracy and grounding. (2) Current VLMs hallucinate frequently even when retrieval is successful; claim-level diagnostics are necessary. (3) Ambiguity is a major failure source: without explicit clarification, models either ignore alternative readings or fuse them incorrectly. (4) Improving evidence selection and utilization—possibly via tighter retrieval–generation coupling or in-the-loop verification—is a promising direction.
@@ A.2 Error Patterns under Ambiguity
    Interpretation conflation: retrieval returns evidence for multiple senses; the generator merges them, producing composite unsupported claims and higher hallucination rates.
@@ Appendix B Retrieve system settings
Table 13: EM/Accuracy performance (%) of each model on the Filtered and Full splits.

### v2

@@ 5.2 End-task Accuracy (short-answer EM/Acc)
Table 2 reports EM/Accuracy across datasets and models. Overall accuracies are modest and vary widely by domain. On the Filtered split, the highest per-dataset scores appear on VisRAG-ArXiv (e.g., up to 63.9 with InternVL3-8B) and VisRAG-Plot (e.g., 51.4 with Pixtral-12B-2409), whereas Visual-RAG and VisRAG-Slide are the most challenging (typically in the 8–24 range across models). WebQA sits in the middle (best around 26.4 on the Filtered split), while Chart-MRAG remains difficult (best Filtered around 22.4).
Comparing Filtered vs. Full, the Full split is generally easier (e.g., Chart-MRAG improves substantially for several models, such as Pixtral-12B-2409 from 14.7 to 29.4), though the effect is not uniform across all datasets and models (e.g., VisRAG-ArXiv sometimes drops for stronger models when moving to Full). These trends confirm that our filtering increases non-triviality and retrieval dependence without relying on a single macro-average.
@@ 5.3 Claim-level Diagnostics (MM-RAGChecker)
Table 3 summarizes claim-level diagnostics on the Filtered split. We observe substantial variation across domains and models rather than a single global rate. Hallucination spans a wide range: it is comparatively low on document-like or chart-heavy sets (e.g., VisRAG-ArXiv around 30–41 for major models; Chart-MRAG can be as low as 24.9 on images for Claude 3.5) but very high on fine-grained visual recognition (Visual-RAG often 74–85). Correspondingly, Faithfulness is highest on structured pages/plots and lowest on open-world visual recognition.
Evidence utilization and cross-modality signals further explain these gaps. On WebQA, Context Precision for images is high (e.g., 89–91), while text exhibits much lower precision (often 28–31), indicating that retrieved passages are frequently unused or unhelpful. Across datasets we also see modality skew: models tend to commit to either text or images instead of reconciling both, which aligns with the elevated hallucination on visually challenging sets and the reduced per-claim F1 on those splits.
Overall, MM-RAGChecker reveals that (i) hallucinations remain frequent on visually demanding queries even when retrieval succeeds, (ii) many retrieved items are not credited as supporting any claim (low context precision, especially on text), and (iii) balancing visual and textual cues is a key bottleneck for long-form grounding.
@@ 5.4 Effect of Retrieval Modes
[…]
@@ 5.5 Prompt Sensitivity Study
We sweep 12 prompt designs (init style, few-shot exemplar count, reasoning pattern). Table 6 lists the best-performing prompt per dataset; Figure 3 visualizes the direct vs. retrieve-then-reason trend. Results in Table 6 are averaged over six models on the Filtered split with fixed seeds and decoding settings. Table 6 reports, for each dataset, the prompt that achieves the highest mean Accuracy/EM across six models (filtered split). We observe clear patterns: Reasoning matters: retrieve_then_reason dominates on text-heavy retrieval tasks (MRAG, WebQA), whereas direct is consistently better on VisRAG-* visual QA splits where context is already focused. Few-shot count is task-dependent: Using three exemplars (ex3) suffices—and sometimes outperforms ex6—for purely visual/numeric QA (plots/slides). In contrast, ex6 helps more on document-style and open-web questions that benefit from pattern coverage. Init style is secondary: Switching between plain and expert seldom changes the rank unless combined with the right reasoning pattern and example count.
@@ 5.7 Effect of Filtering
Filtering not only reduces accuracy but also changes answer style. Filtering increases the average long-answer length from 45 to 60 tokens, yet factuality does not improve: MM-RAGChecker flags a larger fraction of clauses as unsupported on the Filtered split (cf. Table 3). This suggests that harder questions trigger verbosity rather than better grounding.
@@ A.2 Error Patterns under Ambiguity
    Interpretation conflation: retrieval returns evidence for multiple senses; the generator merges them, producing composite neutral claims and higher hallucination rates.
@@ Appendix B Retrieve system settings
Table 13: EM/Accuracy performance (%) of each model on the Filtered and Full splits.All numbers are percentage points; higher is better.